OpenStudy (anonymous):

if two cylinder are similar and the ratio between the lengths of their edges are 2:5. What is the ratio of their volumes?

OpenStudy (anonymous):

Well, volume of a cylinder is \[V= \pi r^2l\] where l is the length of the cylinders edge. The two cylinders are "similar" which I take to mean have a base of equal radii. So the ratio of their volumes is then,\[V _{1}/V _{2}= \pi r^2l _{1}/\pi r^2l _{2}=l _{1}/l _{2}\]Thus the ratio of their volumes is the same as the ratio of their lengths (assuming the radii of the bases are the same). \[V _{1}/V _{2}=l _{1}/l _{2}=2/5\]